# Efficient Discriminative Joint Encoders for Large Scale Vision-Language Reranking

### v1

@@ Abstract
Multimodal retrieval still leans on embedding-based models like CLIP for fast vector search over pre-computed image embeddings. Yet, unlike text retrieval where joint-encoder rerankers are standard, comparable vision–language rerankers are largely absent. We find that seminal joint encoders such as BLIP are severely bottlenecked by an expensive visual feature-extraction stage, preventing practical deployment at scale. Motivated by this bottleneck, we introduce EDJE, an Efficient Discriminative Joint Encoder that precomputes vision tokens offline and compresses them via a lightweight attention-based adapter, so online inference runs only a compact joint encoder over a small set of visual tokens plus the text. EDJE preserves strong retrieval performance while drastically reducing storage and online compute, enabling high-throughput inference. Specifically, EDJE processes 50k image–text pairs/second while requiring 49kB of disk storage per image, matching prior art on Flickr (zero-shot) and COCO (fine-tuned) retrieval. The implementation and checkpoints will be made publicly available shortly.
@@ 1 Introduction
Figure 1: Inference efficiency and retrieval performance. (a) Methods with strong discriminative capabilities are dominated by costly ViT feature extraction, prohibiting their practical use for reranking. (b) EDJE achieves competitive zero-shot retrieval performance with up to $53\times$ faster inference. Its token compression makes storing visual features practical, enabling large-scale retrieval.
@@
A major breakthrough came with the emergence of models that align visual and textual modalities within a shared embedding space, such as CLIP Radford et al. (2021). By enabling efficient similarity search through simple vector comparisons, these models made content-based large-scale retrieval feasible. Beyond retrieval, they have also shown strong generalization to tasks such as zero-shot image classification, inspiring rapid improvements of this paradigm Wang et al. (2022); Zhai et al. (2023); Cherti et al. (2023); Fang et al. (2023a; b); Tschannen et al. (2025).
In parallel, the remarkable success of large language models (LLMs) has driven efforts to integrate vision, enabling instruction-following over multimodal inputs. Early approaches Chen et al. (2020); Li et al. (2021; 2022; 2023) aimed to build foundation vision–language models (VLMs) capable of both generative and discriminative tasks. However, the research community has shifted its interest towards generative-only models, typically by coupling a vision encoder with a pre-trained LLM Liu et al. (2023; 2024); Wang et al. (2024); Gemma (2025). This shift has effectively divided the research community into two main directions: (1) advancing embedding-based models for vision–language alignment, and (2) improving text generation over multimodal inputs — leaving the discriminative potential of joint encoders largely underexplored.
Unlike embedding-based models, joint encoders process both modalities together, allowing richer cross-modal interactions. Prior work Li et al. (2021; 2022; 2023) has shown that such models can significantly improve cross-modal retrieval performance by reranking the top-$k$ candidates retrieved by an embedding model. However, their adoption in practical retrieval pipelines has remained limited; each candidate pair must be evaluated independently, and existing architectures are slow. In particular, these models rely on heavy, high-resolution CLIP-style vision backbones to extract highly expressive image features that poses a severe efficiency bottleneck (Figure˜1(a)). This raises a central question:
@@
To this end, we introduce EDJE, an efficient discriminative joint encoder that allows fine-grained cross-modal interactions without requiring online visual feature extraction. The core idea is to shift visual feature extraction offline: images are encoded once and stored on disk; at query time a compact encoder-only language model jointly processes these with text tokens to produce a reranking score. We further improve scalability by introducing a lightweight token-compression adapter that reduces the number of cached vision tokens. Instead of storing the full sequence produced by the vision backbone, the adapter utilizes a small set of learnable queries that aggregates the most relevant information through cross-attention and projects them to the embedding space of the joint encoder. This compressed representation substantially lowers storage requirements and decreases the number of tokens the joint encoder must process at query time.
Empirically, EDJE consistently improves zero-shot retrieval when paired with a variety of embedding-based models, spanning multiple visual backbones and input resolutions. This demonstrates its modularity as a drop-in reranker that can enhance retrieval pipelines regardless of the underlying embedding model. Moreover, when equipped with a strong visual backbone such as SigLIP2 Tschannen et al. (2025), EDJE surpasses or matches the retrieval performance of prior joint encoders on standard benchmarks (Flickr30k zero-shot; MS-COCO fine-tuned) Plummer et al. (2015); Lin et al. (2014), while operating with substantially greater efficiency (Figure˜1(b)). Finally, we evaluate the robustness of EDJE under compression, quantifying the trade-off between retrieval performance and storage cost as the number of compressed tokens is reduced, and conducting further ablations on reranking pool size and training objectives.
@@
    We introduce EDJE, an efficient discriminative joint encoder that performs fine-grained cross-modal reranking while shifting heavy vision precomputation offline. We further propose a lightweight token-compression adapter that condenses vision features into a compact representation, substantially reducing storage and computation.
@@
    We conduct comprehensive analyses of scalability and robustness, quantifying trade-offs between retrieval performance, storage costs, reranking pool size, and training objectives.
@@ 2 Related Work
Figure 2: Taxonomy of vision–language joint encoders. Left: Cross-attention models integrate modalities through cross-attention layers interleaved with textual self-attention Li et al. (2021; 2022; 2023). Middle: Joint foundation models such as BEiT-3 Wang et al. (2022) employ unified self-attention over native visual and textual tokens, enabling full cross-modal interaction. Right: Modern generative VLMs Liu et al. (2023) combine a pretrained vision encoder with a large language model, tuning the latter to process projected vision tokens as if they originated from text.
@@
The success of CLIP Radford et al. (2021) and ALIGN Jia et al. (2021) in aligning vision and language modalities within a shared embedding space marked a breakthrough in vision–language modeling. By scaling contrastive learning to large architectures and massive image–caption datasets, these models enabled efficient vector similarity search and inspired abundance of follow-up work.
@@
Subsequent research has been directed towards reproducing and extending this paradigm in several directions. For example, LAION-400M Schuhmann et al. (2021) released an open dataset of paired image–caption training data. Other efforts scaled model size and data Xu et al. (2024); Fang et al. (2023b), filtered noisy captions Fang et al. (2023a); Gadre et al. (2023), or generated synthetic ones Li et al. (2022); Nguyen et al. (2023); kokitsi Maninis et al. (2025). Additional work explored alternative loss functions Zhai et al. (2023) or auxiliary objectives to enrich localization Naeem et al. (2024); kokitsi Maninis et al. (2025) and language generation capabilities Wan et al. (2024); Tschannen et al. (2025). Despite their efficiency and scalability, embedding-based approaches compress modalities independently (late interaction), limiting fine-grained cross-modal interactions.
@@
Parallel to contrastive approaches, researchers have pursued models that process modalities jointly. Early systems such as LXMERT and UNITER Tan and Bansal (2019); Chen et al. (2020) relied on region features from R-CNN Ren et al. (2015) combined with text embeddings. The emergence of vision transformers made combining vision and language modalities more straightforward as both modalities are represented as a sequence of tokens. Consequently, some works aimed at creating transformer models capable of processing both images and texts jointly Wang et al. (2021; 2022). Such models require heavy pre-training by masking either text or both text and image tokens. Another line of work introduced cross-attention architectures such as ALBEF, BLIP, BLIP-2, and CoCa Li et al. (2021; 2022; 2023); Yu et al. (2022), which fuse pretrained vision encoders and language models through cross-attention layers. These joint encoders not only unify discriminative and generative modeling, but also consistently outperform embedding-based models on discriminative tasks. In particular, multimodal retrieval performance can be significantly enhanced by reranking embedding-based results with a joint encoder Li et al. (2021; 2022; 2023), echoing common practices in text retrieval where cross-encoders are widely used Chen et al. (2024); Zhang et al. (2024).
@@
A more recent trend is the integration of pretrained vision encoders into large language models (LLMs), yielding generative vision–language models (VLMs). Methods such as LLaVA Liu et al. (2023) introduce a lightweight projection that maps vision tokens into the LLM embedding space, followed by fine-tuning on curated captioning datasets. Variants extend this with parameter-efficient fine-tuning Hu et al. (2022) and supervised fine-tuning Liu et al. (2024); Abdin et al. (2024); Zohar et al. (2025); Microsoft et al. (2025). While this approach makes it simpler to integrate various modalities into highly optimized LLMs Zohar et al. (2025); Microsoft et al. (2025), it often emphasizes instruction-following and generation at the expense of discriminative power.
@@ 3 Towards Efficient Joint Encoders
3.1 On the absence of multimodal rerankers
Existing joint encoders such as BLIP and BLIP-2 Li et al. (2022; 2023) achieve strong performance but rely on visual feature extraction through large backbones like ViT-B/16 (384) and ViT-L/16 (384). This reliance introduces a severe bottleneck: encoding a batch of 64 images requires about 400 ms with ViT-B and nearly 1,400 ms with ViT-L on an A6000 GPU - before any cross-modal interaction even occurs. Specifically, for the BLIP family, the visual feature extraction alone consumes 83% of inference time in the ViT-B case and 93% with ViT-L. In practice, such inference times make it infeasible to use these models for retrieval, where thousands of candidates must be reranked per query. In comparison, the most downloaded text reranker model in HuggingFace [^1] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L6-v2 is based on the MiniLM architecture Wang et al. (2020), has just 22M parameters and processes a similar batch of full-context sequences in under 60 ms, an order of magnitude faster. This gap explains the absence of multimodal rerankers in real-world systems: the cost of extracting visual features alone is prohibitive.
@@ 3.3 Integrating the vision modality
Once visual tokens are computed, the question becomes how best to integrate them with text. Considering how the vision tokens are integrated into "cross-attention" models versus how they are integrated into large vision-language-models, we make an interesting observation: while in "cross attention" models vision tokens are considered in the cross attention layers, large VLMs instead project vision tokens into the language embedding space and concatenate them with text tokens; this allows the cross-modal interaction to be handled entirely by self-attention layers. In our setting, the large language model can be replaced with a compact, efficient language model to meet throughput targets. This yields an architecture with many benefits: (i) Fast: the language model can be as small as MiniLM Wang et al. (2020) or any other efficient language model. (ii) Modular: any ViT-based vision encoder can be paired with any pre-trained language model via a lightweight adapter as a bridge between modalities. (iii) Expressive: modern vision encoders produce highly expressive tokens that capture both semantics and local spatial structure Tschannen et al. (2025). (iv) Data efficient: only the adapter has to be trained from scratch. In the VLM literature it has been observed that the language model and vision encoder require minimal tuning Liu et al. (2023; 2024); Abdin et al. (2024). A high-level view of EDJE is given in Figure˜3(a).
@@ 3.4 Vision-language adapter layer
The adapter projects cached vision tokens into the language embedding space. It has been demonstrated in the VLM literature that even very simple adapters — linear layers Liu et al. (2023) or multi-layer perceptrons (MLPs) Liu et al. (2024) are surprisingly effective despite their parameter count. However, local adapters map each vision token to one language token, limiting flexibility: vision encoders with larger context improve expressiveness but proportionally inflate storage.
@@
This mechanism provides a more flexible way of integrating visual information. Note that it generalizes attention pooling strategies used in embedding models Zhai et al. (2023) and has some connection to the Q-Former layer Li et al. (2023). The token compression layer is depicted in Figure˜3(b).
@@ 4 Effective Discriminative Training
While image–text matching is the central task of interest, it only provides a limited signal for aligning vision–language features and learning a meaningful global representation. Large vision–language models like LLaVA Liu et al. (2023; 2024) achieve this alignment through a pre-training phase in which the model is encouraged to reproduce the caption matching a certain image. Inspired by the pre-training phase of such models and to exploit the bidirectionality of our joint encoder, we employ masked language modeling with aggressive text-only masking. To strengthen the dependence of the [CLS] token on textual inputs, we introduce a projection layer on top of the [CLS] representation and encourage it to recover the text embedding of the underlying embedding model when provided with text-only inputs.
@@ 5.1 Experimental Setup
We train EDJE using a two-phase protocol consisting of pre-training and fine-tuning phases as described in Section˜4. During both phases we freeze the vision encoder and train only the adapter of interest and the language model to process both modalities. We experiment with a variety of vision encoder families at multiple scales and input resolutions, including CLIP Radford et al. (2021), DFN Fang et al. (2023a), MetaCLIP Xu et al. (2024) and SigLIP2 Tschannen et al. (2025). Except for SigLIP2, we use the penultimate-layer hidden states as the vision-encoder output. The language model is fixed to be MiniLM-L12-uncased Wang et al. (2020) in all experiments. To ensure fair comparison, we use the smaller dataset mixture of BLIP for training; the pre-training data is composed of CC12M Changpinyo et al. (2021), CC3M Sharma et al. (2018), SBU Ordonez et al. (2011), Visual Genome Krishna et al. (2016), and COCO Lin et al. (2014), totaling 14M image–caption pairs while fine-tuning only utilizes COCO. Full training hyperparameters are provided in Appendix A.
For evaluation, we follow a two-stage retrieval pipeline: for each query, we first retrieve the top-$k$ candidates using embedding-based retrieval with a CLIP-like model. These candidates are then re-ranked by EDJE, which jointly processes image token embeddings and captions. Unless otherwise stated, we fix the pool-size to $k=10$. We report both text-to-image (T2I) and image-to-text (I2T) performance under $\text{Recall}@\{1,5,10\}$, consistent with prior foundation model benchmarks Li et al. (2021; 2022; 2023); Wang et al. (2022). Since most embedding-based models report other or non-retrieval metrics, we reproduce them using the OpenCLIP framework Ilharco et al. (2021), verifying agreement with their reported numbers before presenting the aforementioned metrics. We evaluate on Flickr30k Plummer et al. (2015) for zero-shot retrieval and on COCO Lin et al. (2014) for fine-tuned retrieval. Following standard practice, we adopt the Karpathy split for COCO and the standard test split of 1,000 images for Flickr30k, each paired with five captions.

### v2

# Efficient Discriminative Joint Encoders for Large Scale Vision-Language Re-ranking

@@ Abstract
Multimodal retrieval still leans on embedding-based models like CLIP for fast vector search over pre-computed image embeddings. Yet, unlike text retrieval where joint-encoder re-rankers are standard, comparable vision–language re-rankers are largely absent. We find that seminal joint encoders such as BLIP are severely bottlenecked by an expensive visual feature-extraction stage, preventing practical deployment at scale. Motivated by this bottleneck, we introduce EDJE, an Efficient Discriminative Joint Encoder that precomputes vision tokens offline and compresses them via a lightweight attention-based adapter, so online inference runs only a compact joint encoder over a small set of visual tokens plus the text. EDJE preserves strong retrieval performance while drastically reducing storage and online compute, enabling high-throughput inference. Specifically, EDJE processes 50k image–text pairs/second while requiring 49kB of disk storage per image, matching prior art on Flickr (zero-shot) and COCO (fine-tuned) retrieval.  [^1] The paper’s code is publicly available at https://github.com/gitanony04-lab/Simple-Efficient-Fusion
@@ 1 Introduction
Figure 1: Inference efficiency and retrieval performance. (a) Methods with strong discriminative capabilities are dominated by costly ViT feature extraction, prohibiting their practical use for re-ranking. (b) EDJE achieves competitive zero-shot retrieval performance with up to $53\times$ faster inference. Its token compression makes storing visual features practical, enabling large-scale retrieval.
@@
A major breakthrough came with the emergence of models that align visual and textual modalities within a shared embedding space, such as CLIP (Radford et al., 2021). By enabling efficient similarity search through simple vector comparisons, these models made content-based large-scale retrieval feasible. Beyond retrieval, they have also shown strong generalization to tasks such as zero-shot image classification, inspiring rapid improvements of this paradigm (Wang et al., 2022; Zhai et al., 2023; Cherti et al., 2023; Fang et al., 2023a; b; Tschannen et al., 2025).
In parallel, the remarkable success of large language models (LLMs) has driven efforts to integrate vision, enabling instruction-following over multimodal inputs. Early approaches (Chen et al., 2020; Li et al., 2021; 2022; 2023) aimed to build foundation vision–language models (VLMs) capable of both generative and discriminative tasks. However, the research community has shifted its interest towards generative-only models, typically by coupling a vision encoder with a pre-trained LLM (Liu et al., 2023; 2024; Wang et al., 2024; Gemma, 2025). This shift has effectively divided the research community into two main directions: (1) advancing embedding-based models for vision–language alignment, and (2) improving text generation over multimodal inputs — leaving the discriminative potential of joint encoders largely underexplored.
Unlike embedding-based models, joint encoders process both modalities together, allowing richer cross-modal interactions. Prior work (Li et al., 2021; 2022; 2023) has shown that such models can significantly improve cross-modal retrieval performance by re-ranking the top-$k$ candidates retrieved by an embedding model. However, their adoption in practical retrieval pipelines has remained limited; each candidate pair must be evaluated independently, and existing architectures are slow. In particular, these models rely on heavy, high-resolution CLIP-style vision backbones to extract highly expressive image features that poses a severe efficiency bottleneck (Figure˜1(a)). This raises a central question:
@@
To this end, we introduce EDJE, an efficient discriminative joint encoder that allows fine-grained cross-modal interactions without requiring online visual feature extraction. The core idea is to shift visual feature extraction offline: images are encoded once and stored on disk; at query time a compact encoder-only language model jointly processes these with text tokens to produce a re-ranking score. We further improve scalability by introducing a lightweight token-compression adapter that reduces the number of cached vision tokens. Instead of storing the full sequence produced by the vision backbone, the adapter utilizes a small set of learnable queries that aggregates the most relevant information through cross-attention and projects them to the embedding space of the joint encoder. This compressed representation substantially lowers storage requirements and decreases the number of tokens the joint encoder must process at query time.
Empirically, EDJE consistently improves zero-shot retrieval when paired with a variety of embedding-based models, spanning multiple visual backbones and input resolutions. This demonstrates its modularity as a drop-in re-ranker that can enhance retrieval pipelines regardless of the underlying embedding model. Moreover, when equipped with a strong visual backbone such as SigLIP2 (Tschannen et al., 2025), EDJE surpasses or matches the retrieval performance of prior joint encoders on standard benchmarks (Flickr30k zero-shot; MS-COCO fine-tuned) (Plummer et al., 2015; Lin et al., 2014), while operating with substantially greater efficiency (Figure˜1(b)). Finally, we evaluate the robustness of EDJE under compression, quantifying the trade-off between retrieval performance and storage cost as the number of compressed tokens is reduced, and conducting further ablations on re-ranking pool size and training objectives.
@@
    We introduce EDJE, an efficient discriminative joint encoder that performs fine-grained cross-modal re-ranking while shifting heavy vision precomputation offline. We further propose a lightweight token-compression adapter that condenses vision features into a compact representation, substantially reducing storage and computation.
@@
    We conduct comprehensive analyses of scalability and robustness, quantifying trade-offs between retrieval performance, storage costs, re-ranking pool size, and training objectives.
@@ 2 Related Work
Figure 2: Taxonomy of vision–language joint encoders. Left: Cross-attention models integrate modalities through cross-attention layers interleaved with textual self-attention (Li et al., 2021; 2022; 2023). Middle: Joint foundation models such as BEiT-3 (Wang et al., 2022) employ unified self-attention over native visual and textual tokens, enabling full cross-modal interaction. Right: Modern generative VLMs (Liu et al., 2023) combine a pretrained vision encoder with a large language model, tuning the latter to process projected vision tokens as if they originated from text.
@@
The success of CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) in aligning vision and language modalities within a shared embedding space marked a breakthrough in vision–language modeling. By scaling contrastive learning to large architectures and massive image–caption datasets, these models enabled efficient vector similarity search and inspired abundance of follow-up work.
@@
Subsequent research has been directed towards reproducing and extending this paradigm in several directions. For example, LAION-400M (Schuhmann et al., 2021) released an open dataset of paired image–caption training data. Other efforts scaled model size and data (Xu et al., 2024; Fang et al., 2023b), filtered noisy captions (Fang et al., 2023a; Gadre et al., 2023), or generated synthetic ones (Li et al., 2022; Nguyen et al., 2023; Maninis et al., 2025). Additional work explored alternative loss functions (Zhai et al., 2023) or auxiliary objectives to enrich localization (Naeem et al., 2024; Maninis et al., 2025) and language generation capabilities (Wan et al., 2024; Tschannen et al., 2025). Despite their efficiency and scalability, embedding-based approaches compress modalities independently (late interaction), limiting fine-grained cross-modal interactions.
Parallel to contrastive approaches, researchers have pursued models that process modalities jointly. Early systems such as LXMERT and UNITER (Tan and Bansal, 2019; Chen et al., 2020) relied on region features from R-CNN (Ren et al., 2015) combined with text embeddings. LightningDOT (Sun et al., 2021) relies on these methods to perform re-ranking with pre-computed region-level representations to enable feasible storage. However, because each region is collapsed into a single vector, such re-ranker behaves much closer to an embedding model rather than a true joint encoder that sees the full image. In practice, this leads to performance that now lags behind modern embedding models such as SigLIP2 (Tschannen et al., 2025).
@@
The emergence of vision transformers made combining vision and language modalities more straightforward as both modalities are represented as a sequence of tokens. Consequently, some works aimed at creating transformer models capable of processing both images and texts jointly (Wang et al., 2021; 2022). Such models require heavy pre-training by masking either text or both text and image tokens. Another line of work introduced cross-attention architectures such as ALBEF, BLIP, BLIP-2, and CoCa (Li et al., 2021; 2022; 2023; Yu et al., 2022), which fuse pretrained vision encoders and language models through cross-attention layers. These joint encoders not only unify discriminative and generative modeling, but also consistently outperform embedding-based models on discriminative tasks. In particular, multimodal retrieval performance can be significantly enhanced by re-ranking embedding-based results with a joint encoder (Li et al., 2021; 2022; 2023), echoing common practices in text retrieval where cross-encoders are widely used (Chen et al., 2024; Zhang et al., 2024).
@@
A more recent trend is the integration of pretrained vision encoders into large language models (LLMs), yielding generative vision–language models (VLMs). Methods such as LLaVA (Liu et al., 2023) introduce a lightweight projection that maps vision tokens into the LLM embedding space, followed by fine-tuning on curated captioning datasets. Variants extend this with parameter-efficient fine-tuning (Hu et al., 2022) and supervised fine-tuning (Liu et al., 2024; Abdin et al., 2024; Zohar et al., 2025; Microsoft et al., 2025). While this approach makes it simpler to integrate various modalities into highly optimized LLMs (Zohar et al., 2025; Microsoft et al., 2025), it often emphasizes instruction-following and generation at the expense of discriminative power.
@@ 3 Towards Efficient Joint Encoders
3.1 On the absence of multimodal re-rankers
Existing joint encoders such as BLIP and BLIP-2 (Li et al., 2022; 2023) achieve strong performance but rely on visual feature extraction through large backbones like ViT-B/16 (384) and ViT-L/16 (384). This reliance introduces a severe bottleneck: encoding a batch of 64 images requires about 400 ms with ViT-B and nearly 1,400 ms with ViT-L on an A6000 GPU - before any cross-modal interaction even occurs. Specifically, for the BLIP family, the visual feature extraction alone consumes 83% of inference time in the ViT-B case and 93% with ViT-L. In practice, such inference times make it infeasible to use these models for retrieval, where thousands of candidates must be re-ranked per query. In comparison, the most downloaded text re-ranker model in HuggingFace [^2] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L6-v2 is based on the MiniLM architecture (Wang et al., 2020), has just 22M parameters and processes a similar batch of full-context sequences in under 60 ms, an order of magnitude faster. This gap explains the absence of multimodal re-rankers in real-world systems: the cost of extracting visual features alone is prohibitive.
@@ 3.3 Integrating the vision modality
Once visual tokens are computed, the question becomes how best to integrate them with text. Considering how the vision tokens are integrated into "cross-attention" models versus how they are integrated into large vision-language-models, we make an interesting observation: while in "cross attention" models vision tokens are considered in the cross attention layers, large VLMs instead project vision tokens into the language embedding space and concatenate them with text tokens; this allows the cross-modal interaction to be handled entirely by self-attention layers. In our setting, the large language model can be replaced with a compact, efficient language model to meet throughput targets. This yields an architecture with many benefits: (i) Fast: the language model can be as small as MiniLM (Wang et al., 2020) or any other efficient language model. (ii) Modular: any ViT-based vision encoder can be paired with any pre-trained language model via a lightweight adapter as a bridge between modalities. (iii) Expressive: modern vision encoders produce highly expressive tokens that capture both semantics and local spatial structure (Tschannen et al., 2025). (iv) Data efficient: only the adapter has to be trained from scratch. In the VLM literature it has been observed that the language model and vision encoder require minimal tuning (Liu et al., 2023; 2024; Abdin et al., 2024). A high-level view of EDJE is given in Figure˜3(a).
@@ 3.4 Vision-language adapter layer
The adapter projects cached vision tokens into the language embedding space. It has been demonstrated in the VLM literature that even very simple adapters — linear layers (Liu et al., 2023) or multi-layer perceptrons (MLPs) (Liu et al., 2024) are surprisingly effective despite their parameter count. However, local adapters map each vision token to one language token, limiting flexibility: vision encoders with larger context improve expressiveness but proportionally inflate storage.
@@
This mechanism provides a more flexible way of integrating visual information. Note that it generalizes attention pooling strategies used in embedding models (Zhai et al., 2023) and has some connection to the Q-Former layer (Li et al., 2023). The token compression layer is depicted in Figure˜3(b).
@@ 4 Effective Discriminative Training
While image–text matching is the central task of interest, it only provides a limited signal for aligning vision–language features and learning a meaningful global representation. Large vision–language models like LLaVA (Liu et al., 2023; 2024) achieve this alignment through a pre-training phase in which the model is encouraged to reproduce the caption matching a certain image. Inspired by the pre-training phase of such models and to exploit the bidirectionality of our joint encoder, we employ masked language modeling with aggressive text-only masking. To strengthen the dependence of the [CLS] token on textual inputs, we introduce a projection layer on top of the [CLS] representation and encourage it to recover the text embedding of the underlying embedding model when provided with text-only inputs.
@@ 5.1 Experimental Setup
We train EDJE using a two-phase protocol consisting of pre-training and fine-tuning phases as described in Section˜4. During both phases we freeze the vision encoder and train only the adapter of interest and the language model to process both modalities. We experiment with a variety of vision encoder families at multiple scales and input resolutions, including CLIP (Radford et al., 2021), DFN (Fang et al., 2023a), MetaCLIP (Xu et al., 2024) and SigLIP2 (Tschannen et al., 2025). Except for SigLIP2, we use the penultimate-layer hidden states as the vision-encoder output. The language model is fixed to be MiniLM-L12-uncased (Wang et al., 2020) in all experiments. To ensure fair comparison, we use the smaller dataset mixture of BLIP for training; the pre-training data is composed of CC12M (Changpinyo et al., 2021), CC3M (Sharma et al., 2018), SBU (Ordonez et al., 2011), Visual Genome (Krishna et al., 2016), and COCO (Lin et al., 2014), totaling 14M image–caption pairs while fine-tuning only utilizes COCO. Full training hyperparameters are provided in Appendix A.
For evaluation, we follow a two-stage retrieval pipeline: for each query, we first retrieve the top-$k$ candidates using embedding-based retrieval with a CLIP-like model. These candidates are then re-ranked by EDJE, which jointly processes image token embeddings and captions. Unless otherwise stated, we fix the pool-size to $k=10$. We report both text-to-image (T2I) and image-to-text (I2T) performance under $\text{Recall}@\{1,5,10\}$, consistent with prior foundation model benchmarks (Li et al., 2021; 2022; 2023; Wang et al., 2022). Since most embedding-based models report other or non-retrieval metrics, we reproduce them using the OpenCLIP framework (Ilharco et al., 2021), verifying agreement with their reported numbers before presenting the aforementioned metrics. We evaluate on Flickr30k (Plummer et al., 2015) for zero-shot retrieval and on COCO (Lin et al., 2014) for fine-tuned retrieval. Following standard practice, we adopt the Karpathy split for COCO and the standard test split of 1,000 images for Flickr30k, each paired with five captions. We additionally evaluate EDJE under a more challenging retrieval setup in which all training images and captions are inserted into the candidate pool, following LightningDOT (Sun et al., 2021). This makes the task considerably more challenging and resemble real-world scenarios. We provide the full details and results in Appendix˜F.